OpenStudy (anonymous):

Solve the equation by completing the square. x^2 + 18x +70 = 0

OpenStudy (anonymous):

Man, it's been a while since I've completed squares. What we're going to do is move the +70 to the other side, so it looks like x^2 + 18x + __ = -70 + __ (these blanks are for when we complete the square!) now, we take half of the coefficient before the x and square it, and then put it into the blank on both sides. so 1/2 of 18 is 9 9^2 = 81 x^2 + 18x + 81 = -70 + 81 x^2 + 18x + 81 = 11 then, we factor the left side (x + 9)(x + 9) = 11 OR (x + 9)^2 = 11 so (x + 9)^2 = 11 take the square of both sides x + 9 = +/- sqrt(11) x = 9 +/- sqrt(11) (because of the +/-, you will have two answers! this is because we took the square root!)

OpenStudy (anonymous):

x^2+18x=-70 x^2+18x+81=-70+81 (x+9)^2=11 Darn It he beat me to it lol

OpenStudy (anonymous):

Sorry! I get excited about completing squares!

OpenStudy (anonymous):

isn't it - 9 +- sqrt 11? because u -9 from x+9

OpenStudy (anonymous):

Ooh, yeah. You're right!

OpenStudy (anonymous):

is x^2 = sqrt of 11 equal to x= +- sqrt of 11? or is it just +-11?

OpenStudy (anonymous):

Man, it's been a while since I've completed squares. What we're going to do is move the +70 to the other side, so it looks like x^2 + 18x + __ = -70 + __ (these blanks are for when we complete the square!) now, we take half of the coefficient before the x and square it, and then put it into the blank on both sides. so 1/2 of 18 is 9 9^2 = 81 x^2 + 18x + 81 = -70 + 81 x^2 + 18x + 81 = 11 then, we factor the left side (x + 9)(x + 9) = 11 OR (x + 9)^2 = 11 so (x + 9)^2 = 11 take the square of both sides x + 9 = +/- sqrt(11) x = 9 +/- sqrt(11) (because of the +/-, you will have two answers! this is because we took the square root!)